OpenStudy (anonymous):

Solve for x: 2/5+3/5x=x+5/10

OpenStudy (anonymous):

fractions are messy so multiply the equation by the common multiple 10 to get rid of them. 10(2/5 + 3/5x) = (x + 5/10)10 distribute the 10 4 + 6x = 10x + 5 isolate the variable from the constants -1 = 4x -1/4 = x